OpenStudy (anonymous):

what's the area of the cube below? Will give medal and fan!!(:

OpenStudy (anonymous):

@ETG311

OpenStudy (candy13106):

d.96

jimthompson5910 (jim_thompson5910):

Use the formula SA = 6s^2 to get SA = 6s^2 SA = 6*4^2 SA = 6*16 SA = 96

OpenStudy (candy13106):

d

OpenStudy (candy13106):

i still think it is d

OpenStudy (anonymous):

thank u evryone! (:

OpenStudy (candy13106):